OpenStudy (anonymous):

∫3x(-e^(-x)) dx Hello, trying to solve this multiplication of integrals and get a wrong answer. Hopefully I can express the ? Lucidly enough that I can understand my mistake. f(x) 3x f'(x) = 3 g(x) and g'(x) = (-e^(-x)) When I multiply f(x) * g(x) it equals 3x(-e^-x) whereas the text and wolfram give the equation as 3x (e^(-x)) Which is then subtracted from g(x) * f'(x) to solve the integral. Why do I get 3x (-e^-x) for f(x) * g(x) rather than the correct 3x (e^(-x)) ?

OpenStudy (mimi_x3):

trying to use integration by parts?

OpenStudy (mimi_x3):

well 3∫e^(-x)xdx u=x dv=e^(-x) du= 1 v= -e^(-x) should be straight forward now.

OpenStudy (anonymous):

So u * v = x * (e^(-x)) not -e^(-x) * x like I figure

OpenStudy (mimi_x3):

well im about to leave so i will leave you the solution for reference :) \[\int 3x(-e^{-x})dx => \int -3xe^{-x}dx => -3\int xe^{-x}dx\] u=x => du=1 dv= e^{-x} => v= -e^(x) \[=>3e^{-x}-3\int e^{-x}dx => 3e^{-x}x+3e^{-x} + C\]

OpenStudy (anonymous):

Ok thanks I will check out your answer.

OpenStudy (raden):

use the table integral to integrate : ∫3x(-e^(-x)) dx =